OpenStudy (anonymous):

how to slove this and make it in to a improper fraction? 2/3 + 1/6 +11/12

OpenStudy (nory):

First we need a common denominator. What is a common multiple of 3, 6, and 12?

OpenStudy (nory):

Some multiples of 3: 3, 6, 9, 12 Some multiples of 6: 6, 12 Some multiples of 12: 12

OpenStudy (anonymous):

common denominator is 12

OpenStudy (nory):

Very good. Now we convert each fraction to a denominator of 12. 11/12 already has a denominator of 12. How about 2/3 and 1/6?

OpenStudy (anonymous):

ok is 2/3 = 8/12 and 1/6= 2/12 is that correct?

OpenStudy (nory):

Very good. Now we can just add them. Add the numerators. The denominator is going to be 12, of course. Then what do you have when you add them? 11/12 + 8/12 + 2/12

OpenStudy (anonymous):

21/12

OpenStudy (nory):

Good. Now do you see a common factor, that we can simplify it with?

OpenStudy (anonymous):

I cant simplify it .

OpenStudy (texaschic101):

your answer 21/12 is an improper fraction

OpenStudy (nory):

@texaschic101 @keshawna It's not in proper form, so shouldn't it be simplified? It is an improper fraction.

OpenStudy (texaschic101):

the questions asks to make an improper fraction....21/12 is an improper fraction

OpenStudy (nory):

You have a point

OpenStudy (texaschic101):

or you can reduce and make it 7/4....that is an improper fraction as well.

OpenStudy (nory):

Yeah